An oil exploration company currently has two active projects, one in Asia and the other in Europe. Let A be the event that the Asian project is successful and 8 be the event that the European project is successful. Suppose that A and B are independent events with A)-0.9 and阀8)-0.5 (a) If the Asian project is not successful, what is the probability that the European project is also not successful? Since the events are independent, then A and B are not independent slce the events are independent, then A and Bare mutually exclusive. Since the events are independent, then A and B are independent. Since the events are not independent, then A and B are mutually exclusive. (b) What is the probability that at least one of the two projects will be successful? 95 bability that only the Asian project is successful? (Round your answer to three decimal (C) Given that at least one of the two projects is successful, what is the pro places.) 0.473 Need Help?Raad Ik to s Tutor Activate Windows
Answer #1

a) P(European project is successful) = 1 - P(European project is successful)

= 1 - 0.5

= 0.5

b) Since the events are independent, then A' and B' are independent

c) P(at least one of the project will be successful) = 1 - P(both projects are not successful)

= 1 - (0.1x0.5)

= 1 - 0.05

= 0.95

d) P(only Asian project is successful | at least one project is successful) = P(only Asian project is successful) P(at least one project is successful)

= P(Asian project is successful and European project is unsuccessful) / 0.95

= 0.9x0.5/0.95

= 0.474